The magnetic field strength at the north pole of a 2.0cm-diameter, 8cm long Alnico magnet is 0.10T. To produce the same field with a solenoid of the same size, carrying a current of 2.1A, how many turns of wire would you need?
Answer #1

B = n*(u) I/l = n*4*10^-7*3.14*2.1*100/8 = 0.10

thus n = 3033 turns

the solenoid is a device that has a large number of turns with the help of thin wire. the wire used is really thin and thus large solnoids can have a huge number of turns. here a length of 8 cm can easily have 3033 turns of the thin wire used in solenoid.